OpenStudy (rational):

Find the form of all positive integers with number of divisors = 10. What is the smallest positive integer ?

OpenStudy (anonymous):

u mean \(\large \tau(n) =10\)

OpenStudy (anonymous):

the smallest :- well i would consider all divisors are multible of 2 so k+1=10 k=9 the smallest would be 2^9

OpenStudy (anonymous):

now hmm idk what do you mean of the FORM :o @rational

OpenStudy (rational):

\[\large (k_1+1)(k_2+1)\cdots = 10\]

OpenStudy (rational):

\[\large (k_1+1)(k_2+1)\cdots = 1\times 2 \times 5\]

OpenStudy (rational):

so we can have 3 prime factors max, right ?

OpenStudy (rational):

\[\large (k_1+1)(k_2+1)(k_3+1) = 1\times 2 \times 5\]

OpenStudy (rational):

sorry i think we can have two prime factors as "1" makes no sense

OpenStudy (anonymous):

make sense , ok keep going

OpenStudy (rational):

\[\large (k_1+1)(k_2+1) = 2 \times 5\]

OpenStudy (anonymous):

wait . ok

OpenStudy (rational):

\(k_1 = 1\) and \(k_2 = 4\) or \(k_1 = 4\) and \(k_2 = 1\)

OpenStudy (rational):

wait, ur case should be considered as well : \[k_1 + 1 = 10\] \(k_1 = 9\)

OpenStudy (rational):

so the minimum number would be : min{2^9, 2*3^4, 2^4*3}

OpenStudy (anonymous):

my case only gives min

OpenStudy (rational):

it doesnt

OpenStudy (anonymous):

oh :o

OpenStudy (rational):

min is 2^4*3 = 48

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ok what about the form ?

OpenStudy (rational):

try another easy proof ? :) http://prntscr.com/4fludo

OpenStudy (rational):

form is simply \(\large n = p^9\) or \(\large n = p_1p_2^4\)

OpenStudy (anonymous):

ok keep going with part b

OpenStudy (anonymous):

part b >.< not easy to me

OpenStudy (rational):

since \( n \lt \sigma(n)\) we only need to test \(n \lt 10\)

OpenStudy (rational):

\(\sigma(2) = 1+2 = 3\)

OpenStudy (rational):

\(\sigma(3) = 1+3 = 4\) \(\sigma(4) = 1+2 + 4 = 7\) \(\sigma(5) = 1+5 = 6\)

OpenStudy (anonymous):

then ?

OpenStudy (rational):

\(\sigma(6) = 1+2+3+6 = 12\) \(\sigma(7) = 1+7=8\) \(\sigma(8) = 1+2+4+8 = 15\) \(\sigma(9) = 1+3+9= 13\) for \(n \gt 9\), since the \(\sigma(n) \gt 10\), there will not be any numbers with \(\sigma(n)=10\)

OpenStudy (anonymous):

ahaa ! cool >.<

OpenStudy (rational):

one more http://prntscr.com/4flwjz

OpenStudy (anonymous):

shoul i use \(\sigma (m^2)=\sigma (m)\sigma (m)\)

OpenStudy (rational):

thats not true as \(\large \gcd(m,m) \ne 1\)

OpenStudy (anonymous):

>.<

OpenStudy (rational):

\(\large \sigma (4 . 2) = 1+2+4+8 = 15 \color{red}{\ne} 21 = \sigma(4)*\sigma(2)\)

OpenStudy (rational):

multiplicative works only when gcd is 1

OpenStudy (anonymous):

roger !

OpenStudy (anonymous):

ok so k could be any number , or i can chose w.e i want ?

OpenStudy (rational):

yeah say \(k = 3\) ?

OpenStudy (anonymous):

got it !

OpenStudy (rational):

how ?

OpenStudy (anonymous):

there is infinitly many number combination of 4k, 5k \(\sigma(m^2)=\sigma(4^2 k^2)= \sigma(4^2) \sigma (k^2)=\sigma(5^2)\sigma(k^2)=\sigma (5^2k^2)=\sigma(n^2)\)

OpenStudy (rational):

got it :)

OpenStudy (rational):

thats the reason we need gcd(5,4,k) = 1 is it

OpenStudy (rational):

so that we can do \(\large \sigma(4^2k^2) = \sigma(4^2)*\sigma(k^2)\)

OpenStudy (anonymous):

yeah , we can chose other forms like bk,ak but we need \(\sigma( a^2) =\sigma (b^2))

OpenStudy (anonymous):

\(\sigma( a^2) =\sigma (b^2)\)

OpenStudy (rational):

and \(\large \sigma(4^2) = 31 = \sigma(5^2)\)

OpenStudy (rational):

understood :)

OpenStudy (anonymous):

yosh